For what value(s) of the parameter m does the equation -2x^2 + mx = 2 have one solution only?
A) 0
B) -2 , 2
C) -1 , 1
D) 16
E) -4 , 4

Answer by mananth(16949)

-2x^2 + mx = 2 have one solution only?
..
rewrite the equation
2x^2-mx=-2
2x^2-mx+2=0
/2
x^2-mx/2 +1=0
s0 m= +/-4
...
m.ananth@hotmail.ca